OpenStudy (anonymous):

At a restaurant, 5 of every 12 customers had a salad with their meal. Thirty-five customers had a salad with their meal. How many customers were there altogether?

OpenStudy (anonymous):

We need to cross multiply again!

OpenStudy (anonymous):

5/12 = 35/x What do you get after cross multiplying?

OpenStudy (anonymous):

five divided by twelve?

OpenStudy (anonymous):

wat?

OpenStudy (anonymous):

12/4=2.4 2.4x35=84?

OpenStudy (anonymous):

5 customers out of 12 customers have a salad 35 customers out of x customers have a salad. We are solving for 'x'

OpenStudy (anonymous):

84.

OpenStudy (anonymous):

Correct!

OpenStudy (anonymous):

12/5=2.4 2.4x35=84

OpenStudy (anonymous):

ty

OpenStudy (anonymous):

My pleasure!`

OpenStudy (anonymous):

My pleasure!*